• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Eric Voutier

    Log entry time 00:07:33 on December 13,1999

    Entry number 30911

    keyword=Shift summary

    16:00 Shift change-over; we are starting the study of the beam displacement effect
    for the calibration of the spot++ program of Bob.

    Note: the acquisition dead-time is close to 50%, Michael is investigating the
    problem... no real idea about this.

    run 1232 beam +2mm right
    run 1233 beam -2mm left
    run 1234 beam +2mm up
    run 1235 beam -2mm down

    17:15 Q2 in hadron is negative polarity??.... we perform again the kapton and carbon measurement in (e,e'p) after bringing back the quad at its right setting.

    Note: as observed by Calarco, it seems that the ramp-down speed of the power supply of the magnets is too fast when set by the selected momentum. The proposal
    is to go up to 1200 A, then wait and after go to 1000 A and finally fixed the current
    by the chosen mometum.... we follow it and the magnet did not trip.

    17:50 Start carbon data acquisition, the coincidence event rate is about 10Hz

    run 1236 ps1/2/3/4/5/8 20/20/30/30/1/65535 ~10 k (T5)
    run 1237 ps1/2/3/4/5/8 30/30/50/50/1/65535 ~13 k
    run 1238 ps1/2/3/4/5/8 idem ~13 k
    run 1239 ps1/2/3/4/5/8 idem
    run 1240 ps1/2/3/4/5/8 idem
    run 1241 ps1/2/3/4/5/8 idem
    run 1242 ps1/2/3/4/5/8 idem
    run 1243 ps1/2/3/4/5/8 idem

    This should give in the overall about 100 k T5 triggers.

    19:30 Stop for a quater of an hour to let Hall B determine the origin of their 50pA
    current leak... it turns out that they are sensitive to Hall C and dont see Hall A at

    21:00 Move to the 10 cm dummy target for data taking at quasielastic kinematics
    run 1245 ps1/2/3/4/5/8 30/30/50/50/1/65535 (5uA)
    run 1246 idem
    run 1247 idem
    run 1248 ps1/2/3/4/5/8 100/100/100/100/1/6535 (30uA)
    run 1249

    22:30 Move the spectrometer tyo kinematics #17 settings

    We get troubles with the NMR probes of the dipole...